OpenStudy (anonymous):

An equation is shown 7x+y=5 Part A: Explain how you will show all of the solutions that satisfy this equation.

OpenStudy (anonymous):

You'd graph it!

OpenStudy (anonymous):

But it's unreasonable to show every solution, because there is a solution to this equation from negative infinity to infinity.

OpenStudy (mathmale):

All you have to do is to graph this straight line. That line shows all of the solutions that satisfy this equation. 2 methods of graphing: 1) "Intercept" method: if we let x=0, 0+y=5, so y=5 and the point is (0,5); if we let y=0, 7x+0=5, or x=5/7, and another point to graph is (5/7, 0). 2) slope-intercept formula method; Solve 7x+y=5 for y. Answer: y=-7/x + 5. Choose several x values and find the associated y values. Graph these points and then draw a straight line through them.

OpenStudy (anonymous):

THEN IT ASKS determin three different solution for the equation

OpenStudy (anonymous):

determine*

OpenStudy (mathmale):

If you followed my Example 2, you have already determined 2 or 3 different solutions to this equation. For example, if you let x=0, then y = 7x + 5 becomes y = 7(0)+5 = 5 and the solution is just (0,5). Let x=1 and find another solution. Let x =3 and find another solution.

OpenStudy (anonymous):

for 1 I got (1, -3)